Circular No. 6034

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

PERIODIC COMET SHOEMAKER-LEVY 9 (1993e) T. Encrenaz, R. Schulz, J. A. Stuewe and G. Wiedemann report: “We

monitored Jupiter in the 3.0–3.5-μm range shortly after the impact of fragment H using the IRSPEC imaging spectrometer on the 3.5-m New Technology Telescope at the European Southern Observatory. Our observations (resolving power 1700) started on July 18.824 UT (13 min after impact) at a rate of one spectrum per minute. Very strong emission was detected over the whole spectral range from multiplets of the CH_{4} ν_{3} band (J = 14 to J = 18). A spatial average over the impact region shows that the signal intensity decreased exponentially with a timescale of ∼ 5 min. The IRSPEC spectrometer, its 4′′.4 slit aligned along the parallel (latitude −44^{o}) of the impacts, allows a spatial analysis (spatial resolution 2′′.2) in this direction. At the beginning of the sequence, the CH_{4} emission extended over ∼ 10′′. The slope of the spectrum shows drastic variations along the slit, corresponding to a large range of rotational temperatures. In the first image, the peak of intensity (at the center of the emission) corresponds to a rotational temperature of ∼ 700 K. On the leading side, the intensity is much weaker, but its rotational temperature reaches several thousand degrees. Both the hot and the central regions cooled gradually to temperatures of ∼ 1000 and 500 K, respectively, 15 min after the first image.”

Reports received via the SL9 message center indicate that there were no obvious detections of the impacts of fragments T = 4 (Calar Alto Observatory, South African Astronomical Observatory), U = 3 (Palomar Observatory, Teide Observatory, Swedish Solar Telescope at La Palma, University of Massachusetts Whately Observatory, McDonald Observatory) or V = 2 (McDonald, Palomar, Steward Observatory, Siding Spring Observatory).

The impact of fragment W = 1 was recorded by SPIREX (cf. IAUC 6026) on July 22.343 UT, a preliminary estimate indicating that the site was as bright as that of fragment E at 2.36 μm. CASPIR (cf. IAUC 6024) detected a slight brightening at 2.34 μm on July 22.337 that faded briefly by 22.340, then strongly reappeared, saturating the detector on 22.342 as the W site came into view; narrow-band imaging beginning on 22.345 showed strong emission at 4.78 μm. IRIS K-grism cubes (cf. IAUC 6024) showed the W impact as a distinct, bright, pointlike source slightly south of the K site on July 22.342; the W feature brightened dramatically, saturating the detector on 22.344, when it was > 200 times brighter than the south polar hood and ∼ 4.5 times brighter than Ganymede at 2.34 μm.

A. C. Beresford, Myrtle Bank, South Australia, observed the W impact site visually as it crossed Jupiter’s central meridian on July 22.436 UT.

1994 July 22 Brian G. Marsden